A unit vector perpendicular to plane determined by the points (1,-1,2) (2,0,-1) and (0,2,1) is

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavui

poinys are given in a plane

so two vactor in that plane can be obtain

a= (2-1) i +( 0+1)j +(-1-2)k

   = i +j -3k

and   b= (0-1)i + (2+1) j +(1-2)k

          =-i +3j -k

now a vector perpendicular to plane is   r =a X b

for unit vector

          r= (a X b)/|a X b|

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin